# Mental Disorders Detection in the Era of Large Language Models

### v1

@@ Appendix B Full Results for Models
| DE | Gemma2 2B IT | 5-shot MMLU | 13.33 | 80.0 | 22.86 | 86.36 | 19.59 | 31.93 | 27.39 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 13.33 | 100.0 | 23.53 | 100.0 | 22.0 | 36.07 | 29.8 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 73.33 | 92.96 | 81.99 | 77.27 | 41.46 | 53.97 | 67.98 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 53.33 | 92.31 | 67.61 | 81.82 | 30.0 | 43.9 | 55.75 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 25.56 | 88.46 | 39.66 | 86.36 | 22.09 | 35.19 | 37.42 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 38.89 | 92.11 | 54.69 | 86.36 | 25.68 | 39.58 | 47.14 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 18.89 | 85.0 | 30.91 | 86.36 | 20.65 | 33.33 | 32.12 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 35.56 | 88.89 | 50.79 | 81.82 | 23.68 | 36.73 | 43.76 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 14.44 | 72.22 | 24.07 | 77.27 | 18.09 | 29.31 | 26.69 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 19.64 | 32.84 | 16.42 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 82.22 | 87.06 | 84.57 | 50.0 | 40.74 | 44.9 | 64.73 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 5.56 | 71.43 | 10.31 | 90.91 | 19.05 | 31.5 | 20.9 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 92.22 | 85.57 | 88.77 | 36.36 | 53.33 | 43.24 | 66.01 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 97.78 | 80.0 | 88.0 | 0.0 | 0.0 | 0.0 | 44.0 | |||||||||
| DSM | Gemma2 2B IT | 5-shot MMLU | 25.93 | 70.0 | 37.84 | 83.33 | 42.86 | 56.6 | 47.22 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 44.44 | 60.0 | 51.06 | 55.56 | 40.0 | 46.51 | 48.79 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 85.19 | 65.71 | 74.19 | 33.33 | 60.0 | 42.86 | 58.53 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 59.26 | 64.0 | 61.54 | 50.0 | 45.0 | 47.37 | 54.45 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 44.44 | 66.67 | 53.33 | 66.67 | 44.44 | 53.33 | 53.33 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 62.96 | 62.96 | 62.96 | 44.44 | 44.44 | 44.44 | 53.7 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 3.7 | 100.0 | 7.14 | 100.0 | 40.91 | 58.06 | 32.6 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 51.85 | 63.64 | 57.14 | 55.56 | 43.48 | 48.78 | 52.96 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 100.0 | 60.0 | 75.0 | 0.0 | 0.0 | 0.0 | 37.5 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 83.33 | 35.71 | 50.0 | 25.0 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 81.48 | 68.75 | 74.58 | 44.44 | 61.54 | 51.61 | 63.09 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 51.85 | 60.87 | 56.0 | 50.0 | 40.91 | 45.0 | 50.5 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 92.59 | 64.1 | 75.76 | 22.22 | 66.67 | 33.33 | 54.55 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 60.0 | 75.0 | 0.0 | 0.0 | 0.0 | 37.5 | |||||||||
| AL | Gemma2 2B IT | 5-shot MMLU | 4.55 | 50.0 | 8.33 | 94.74 | 46.15 | 62.07 | 35.2 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 46.34 | 63.33 | 31.67 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 45.45 | 66.67 | 54.05 | 73.68 | 53.85 | 62.22 | 58.14 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 50.0 | 64.71 | 56.41 | 68.42 | 54.17 | 60.47 | 58.44 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 27.27 | 60.0 | 37.5 | 78.95 | 48.39 | 60.0 | 48.75 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 46.34 | 63.33 | 31.67 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 31.82 | 63.64 | 42.42 | 78.95 | 50.0 | 61.22 | 51.82 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 86.36 | 55.88 | 67.86 | 21.05 | 57.14 | 30.77 | 49.31 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 4.55 | 33.33 | 8.0 | 89.47 | 44.74 | 59.65 | 33.82 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 77.27 | 53.12 | 62.96 | 21.05 | 44.44 | 28.57 | 45.77 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 53.66 | 69.84 | 0.0 | 0.0 | 0.0 | 34.92 | |||||||||
| AD | Gemma2 2B IT | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 10.0 | 50.0 | 16.67 | 88.89 | 47.06 | 61.54 | 39.1 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 15.0 | 50.0 | 23.08 | 83.33 | 46.88 | 60.0 | 41.54 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 35.0 | 77.78 | 48.28 | 88.89 | 55.17 | 68.09 | 58.18 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 70.0 | 73.68 | 71.79 | 72.22 | 68.42 | 70.27 | 71.03 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 20.0 | 80.0 | 32.0 | 94.44 | 51.52 | 66.67 | 49.33 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 85.0 | 53.12 | 65.38 | 16.67 | 50.0 | 25.0 | 45.19 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 25.0 | 83.33 | 38.46 | 94.44 | 53.12 | 68.0 | 53.23 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 10.0 | 33.33 | 15.38 | 77.78 | 43.75 | 56.0 | 35.69 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 52.63 | 68.97 | 0.0 | 0.0 | 0.0 | 34.48 | |||||||||
| AC | Gemma2 2B IT | 5-shot MMLU | 17.78 | 72.73 | 28.57 | 92.11 | 48.61 | 63.64 | 46.1 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 17.78 | 72.73 | 28.57 | 92.11 | 48.61 | 63.64 | 46.1 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 88.89 | 57.14 | 69.57 | 21.05 | 61.54 | 31.37 | 50.47 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 35.56 | 61.54 | 45.07 | 73.68 | 49.12 | 58.95 | 52.01 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 31.11 | 66.67 | 42.42 | 81.58 | 50.0 | 62.0 | 52.21 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 48.89 | 59.46 | 53.66 | 60.53 | 50.0 | 54.76 | 54.21 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 13.33 | 100.0 | 23.53 | 100.0 | 49.35 | 66.09 | 44.81 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 42.22 | 70.37 | 52.78 | 78.95 | 53.57 | 63.83 | 58.3 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 100.0 | 54.22 | 70.31 | 0.0 | 0.0 | 0.0 | 35.16 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 40.0 | 64.29 | 49.32 | 73.68 | 50.91 | 60.22 | 54.77 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 17.78 | 53.33 | 26.67 | 81.58 | 45.59 | 58.49 | 42.58 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 51.11 | 57.5 | 54.12 | 55.26 | 48.84 | 51.85 | 52.98 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 97.78 | 53.66 | 69.29 | 0.0 | 0.0 | 0.0 | 34.65 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 97.78 | 53.66 | 69.29 | 0.0 | 0.0 | 0.0 | 34.65 |

### v3

@@ Appendix B Full Results for Models
| DE | Gemma2 2B IT | 5-shot MMLU | 80.00 0.00 | 13.33 0.00 | 22.86 0.00 | 19.59 0.00 | 86.36 0.00 | 31.93 0.00 | 27.39 0.00 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 100.00 0.00 | 13.33 0.00 | 23.53 0.00 | 22.00 0.00 | 100.00 0.00 | 36.07 0.00 | 29.80 0.00 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 93.11 0.08 | 75.11 0.89 | 83.15 0.58 | 43.16 0.85 | 77.27 0.00 | 55.38 0.71 | 69.26 0.64 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 92.31 0.00 | 53.33 0.00 | 67.61 0.00 | 30.00 0.00 | 81.82 0.00 | 43.90 0.00 | 55.75 0.00 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 88.80 0.17 | 26.44 0.44 | 40.75 0.55 | 22.30 0.10 | 86.36 0.00 | 35.45 0.13 | 38.10 0.34 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 92.42 0.16 | 40.67 0.89 | 56.48 0.89 | 26.25 0.29 | 86.36 0.00 | 40.26 0.34 | 48.37 0.62 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 85.00 0.00 | 18.89 0.00 | 30.91 0.00 | 20.65 0.00 | 86.36 0.00 | 33.33 0.00 | 32.12 0.00 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 88.37 0.26 | 33.78 0.89 | 48.87 0.96 | 23.20 0.24 | 81.82 0.00 | 36.15 0.29 | 42.51 0.63 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 74.44 1.11 | 16.22 0.89 | 26.63 1.28 | 18.40 0.16 | 77.27 0.00 | 29.72 0.21 | 28.18 0.74 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 19.64 0.00 | 100.00 0.00 | 32.84 0.00 | 16.42 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 87.06 0.00 | 82.22 0.00 | 84.57 0.00 | 40.74 0.00 | 50.00 0.00 | 44.90 0.00 | 64.73 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 71.43 0.00 | 5.56 0.00 | 10.31 0.00 | 19.05 0.00 | 90.91 0.00 | 31.50 0.00 | 20.90 0.00 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 85.45 0.06 | 91.33 0.44 | 88.29 0.24 | 50.67 1.33 | 36.36 0.00 | 42.33 0.46 | 65.31 0.35 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 80.14 0.07 | 98.67 0.44 | 88.45 0.22 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 44.22 0.11 | |||||||||
| DSM | Gemma2 2B IT | 5-shot MMLU | 71.14 0.57 | 20.00 2.96 | 31.10 3.37 | 42.26 0.30 | 87.78 2.22 | 57.04 0.22 | 44.07 1.58 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 60.00 0.00 | 44.44 0.00 | 51.06 0.00 | 40.00 0.00 | 55.56 0.00 | 46.51 0.00 | 48.79 0.00 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 64.91 0.40 | 82.22 1.48 | 72.54 0.82 | 55.64 2.18 | 33.33 0.00 | 41.67 0.59 | 57.11 0.71 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 64.00 0.00 | 59.26 0.00 | 61.54 0.00 | 45.00 0.00 | 50.00 0.00 | 47.37 0.00 | 54.45 0.00 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 66.67 0.00 | 44.44 0.00 | 53.33 0.00 | 44.44 0.00 | 66.67 0.00 | 53.33 0.00 | 53.33 0.00 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 62.96 0.00 | 62.96 0.00 | 62.96 0.00 | 44.44 0.00 | 44.44 0.00 | 44.44 0.00 | 53.70 0.00 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 100.00 0.00 | 3.70 0.00 | 7.14 0.00 | 40.91 0.00 | 100.00 0.00 | 58.06 0.00 | 32.60 0.00 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 63.64 0.00 | 51.85 0.00 | 57.14 0.00 | 43.48 0.00 | 55.56 0.00 | 48.78 0.00 | 52.96 0.00 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 60.00 0.00 | 100.00 0.00 | 75.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 37.50 0.00 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 35.71 0.00 | 83.33 0.00 | 50.00 0.00 | 25.00 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 68.75 0.00 | 81.48 0.00 | 74.58 0.00 | 61.54 0.00 | 44.44 0.00 | 51.61 0.00 | 63.09 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 60.87 0.00 | 51.85 0.00 | 56.00 0.00 | 40.91 0.00 | 50.00 0.00 | 45.00 0.00 | 50.50 0.00 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 62.34 0.88 | 95.56 1.48 | 75.44 0.16 | 66.67 0.00 | 13.33 4.44 | 21.90 5.71 | 48.67 2.94 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 60.00 0.00 | 100.00 0.00 | 75.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 37.50 0.00 | |||||||||
| AL | Gemma2 2B IT | 5-shot MMLU | 50.00 0.00 | 4.55 0.00 | 8.33 0.00 | 46.15 0.00 | 94.74 0.00 | 62.07 0.00 | 35.20 0.00 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 46.34 0.00 | 100.00 0.00 | 63.33 0.00 | 31.67 0.00 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 66.67 0.00 | 45.45 0.00 | 54.05 0.00 | 53.85 0.00 | 73.68 0.00 | 62.22 0.00 | 58.14 0.00 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 64.71 0.00 | 50.00 0.00 | 56.41 0.00 | 54.17 0.00 | 68.42 0.00 | 60.47 0.00 | 58.44 0.00 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 60.00 0.00 | 27.27 0.00 | 37.50 0.00 | 48.39 0.00 | 78.95 0.00 | 60.00 0.00 | 48.75 0.00 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 61.03 0.52 | 85.45 1.82 | 71.20 0.99 | 68.73 2.55 | 36.84 0.00 | 47.95 0.64 | 59.58 0.82 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 46.34 0.00 | 100.00 0.00 | 63.33 0.00 | 31.67 0.00 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 63.64 0.00 | 31.82 0.00 | 42.42 0.00 | 50.00 0.00 | 78.95 0.00 | 61.22 0.00 | 51.82 0.00 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 55.88 0.00 | 86.36 0.00 | 67.86 0.00 | 57.14 0.00 | 21.05 0.00 | 30.77 0.00 | 49.31 0.00 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 45.00 0.00 | 94.74 0.00 | 61.02 0.00 | 30.51 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 60.00 0.00 | 81.82 0.00 | 69.23 0.00 | 63.64 0.00 | 36.84 0.00 | 46.67 0.00 | 57.95 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 33.33 0.00 | 4.55 0.00 | 8.00 0.00 | 44.74 0.00 | 89.47 0.00 | 59.65 0.00 | 33.82 0.00 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 51.92 0.60 | 73.64 1.82 | 60.89 1.03 | 40.89 1.78 | 21.05 0.00 | 27.78 0.39 | 44.34 0.71 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 53.66 0.00 | 100.00 0.00 | 69.84 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 34.92 0.00 | |||||||||
| AD | Gemma2 2B IT | 5-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 47.37 0.00 | 100.00 0.00 | 64.29 0.00 | 32.14 0.00 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 50.00 0.00 | 10.00 0.00 | 16.67 0.00 | 47.06 0.00 | 88.89 0.00 | 61.54 0.00 | 39.10 0.00 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 55.71 2.86 | 19.00 2.00 | 28.32 2.62 | 48.08 0.60 | 83.33 0.00 | 60.98 0.49 | 44.65 1.56 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 77.78 0.00 | 35.00 0.00 | 48.28 0.00 | 55.17 0.00 | 88.89 0.00 | 68.09 0.00 | 58.18 0.00 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 70.74 1.47 | 70.00 0.00 | 70.36 0.72 | 67.02 0.70 | 67.78 2.22 | 67.39 1.44 | 68.87 1.08 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 47.37 0.00 | 100.00 0.00 | 64.29 0.00 | 32.14 0.00 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 80.00 0.00 | 20.00 0.00 | 32.00 0.00 | 51.52 0.00 | 94.44 0.00 | 66.67 0.00 | 49.33 0.00 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 55.62 1.25 | 89.00 2.00 | 68.46 1.54 | 63.33 6.67 | 21.11 2.22 | 31.67 3.33 | 50.06 2.44 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 13.33 26.67 | 2.00 4.00 | 3.48 6.96 | 46.47 1.05 | 94.44 0.00 | 62.28 0.93 | 32.88 3.94 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 66.67 0.00 | 10.00 0.00 | 17.39 0.00 | 48.57 0.00 | 94.44 0.00 | 64.15 0.00 | 40.77 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 83.33 0.00 | 25.00 0.00 | 38.46 0.00 | 53.12 0.00 | 94.44 0.00 | 68.00 0.00 | 53.23 0.00 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 38.67 2.67 | 10.00 0.00 | 15.88 0.25 | 45.11 0.68 | 82.22 2.22 | 58.26 1.13 | 37.07 0.69 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 52.63 0.00 | 100.00 0.00 | 68.97 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 34.48 0.00 | |||||||||
| AC | Gemma2 2B IT | 5-shot MMLU | 74.55 0.91 | 19.56 0.89 | 30.98 1.20 | 49.16 0.27 | 92.11 0.00 | 64.10 0.23 | 47.54 0.72 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 72.73 0.00 | 17.78 0.00 | 28.57 0.00 | 48.61 0.00 | 92.11 0.00 | 63.64 0.00 | 46.10 0.00 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 60.46 1.66 | 51.56 18.67 | 53.91 7.83 | 52.31 4.62 | 58.95 18.95 | 52.50 10.56 | 53.20 1.37 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 58.46 1.54 | 33.78 0.89 | 42.82 1.13 | 47.72 0.70 | 71.58 1.05 | 57.26 0.84 | 50.04 0.98 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 71.52 2.42 | 34.67 1.78 | 46.69 2.13 | 51.97 0.98 | 83.68 1.05 | 64.12 1.06 | 55.41 1.60 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 59.89 0.22 | 52.44 1.78 | 55.91 1.12 | 50.93 0.47 | 58.42 1.05 | 54.41 0.18 | 55.16 0.47 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 100.00 0.00 | 13.33 0.00 | 23.53 0.00 | 49.35 0.00 | 100.00 0.00 | 66.09 0.00 | 44.81 0.00 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 67.41 1.48 | 40.44 0.89 | 50.56 1.11 | 52.14 0.71 | 76.84 1.05 | 62.13 0.85 | 56.34 0.98 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 54.22 0.00 | 100.00 0.00 | 70.31 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 35.16 0.00 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 69.15 2.43 | 41.78 0.89 | 52.09 1.39 | 53.04 1.06 | 77.89 2.11 | 63.11 1.45 | 57.60 1.42 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 90.67 18.67 | 8.89 4.44 | 15.33 5.67 | 47.12 0.76 | 96.32 7.37 | 63.22 2.37 | 39.28 1.65 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 57.50 0.00 | 51.11 0.00 | 54.12 0.00 | 48.84 0.00 | 55.26 0.00 | 51.85 0.00 | 52.98 0.00 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 55.29 0.82 | 97.78 0.00 | 70.63 0.67 | 60.00 30.00 | 6.32 3.16 | 11.43 5.71 | 41.03 3.19 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 53.66 0.00 | 97.78 0.00 | 69.29 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 34.65 0.00 |